OpenStudy (anonymous):

Based on the lesson, which of the following would be the best approach for solving this system by substitution? 5x = y + 6 2x - 3y = 4 Solve the first equation for x. Solve the first equation for y. Solve the second equation for x. Solve the second equation for y.

OpenStudy (shootingstar28):

Okay, so it says we have to use SUBSTITUTION. That means we have to solve for a variable in one of either equations. Which equation would be the easiest to work with and use in this case?

OpenStudy (shootingstar28):

Here are our two equations: 5x = y + 6 2x - 3y = 4 As you can see now, that the first equation already has a variable set, y. Now you should solve for y for the first equation.

OpenStudy (shootingstar28):

The second approach would have to be the simplest, the reason why is because the y is easy to isolate. y=5x-6

OpenStudy (shootingstar28):

So pretty much that can be substituted in the second equation: 2x -3(5x-6)=4 2x-15x + 18 = 4 -13x=-18+4 -13x = -14 x=14/13

OpenStudy (anonymous):

thank you:)

OpenStudy (shootingstar28):

Welcome! :3